I need some help rewriting this expression using the distributive property:
8x-40y
Thanks in advance for your help

Answer by praseenakos@yahoo.com(507)

QUESTION:
rewriting this expression using the distributive property:
8x-40y
ANSWER:

The Distributive Property is an algebra property which is used to multiply a single term and two or more terms inside a set of parentheses.

Take a look at the problem below.


2(3 + 6) = 2*3 + 2* 6



Your Question:

8x-40y



Here take out common terms from 8x and 40y


8x-40y = 8x - 5 * 8y


Here 8 is common in both the terms


==> 8x-40y = 8 ( 5 - 5y )


which is the required answer.



Hope you understood.
Regards,
praseenakos@yahoo.co.in